# Port-LLM: A Port Prediction Method for Fluid Antenna based on Large Language Models

### v1

@@ V Conclusions
In this paper, the FA is employed to mitigate the mobility-induced challenges in communication systems. Notably, leveraging the power of the LLMs, we present a moving port prediction model for the FA, designated as Port-LLM. To the best of our current knowledge, this study represents the first application of neural networks to moving port prediction within the domain of FA. By repositioning the FA to the port forecasted by our proposed model, it becomes feasible to maintain an approximately invariant channel state information as the UE moves. The port prediction approach based on our Port-LLM model primarily consists of two sequential steps: the first step involves predicting the channel table, while the second step pertains to port prediction. Furthermore, during the training phase of the Port-LLM model, we incorporated LoRA fine-tuning, which substantially curtailed the number of parameters necessitating retraining. We introduced the warm-up-aided cosine LR technique to improve the prediction accuracy and convergence speed of our model. Simulation findings indicate that our proposed model exhibits pronounced performance enhancements compared to traditional techniques, particularly in medium and high-speed scenarios. Concurrently, our model demonstrates strong robustness under diverse BS antenna configurations and varying UE movement velocities.

### v2

@@ V Conclusions
In this paper, the FA is employed to mitigate mobility-induced challenges in communication systems. In particular, leveraging the powerful modeling capabilities of LLMs, we propose a Port-LLM model based on LoRA fine-tuning. Building on this, we also propose a Prompt-Port-LLM model based on prompt fine-tuning to further utilize the exceptional NLP capabilities of LLMs. By repositioning the FA to the port predicted by our proposed models, it becomes feasible to maintain an approximately invariant channel state information as the UE moves. To effectively address the discrepancies between the wireless communication data format and the input format of the pre-trained LLM, we specially design the data processing module, input embedding module and output projection module. Moreover, in our proposed Prompt-Port-LLM model, we design specialized dynamic prompts and a dedicated prompt encoder module, thereby further enhancing the model’s prediction accuracy. Simulation results show that both of our proposed models exhibit significant performance improvement compared to the traditional technical solutions, especially in medium and high-speed scenarios. In addition, our proposed models show good robustness under different BS-side antenna configurations and different UE movement speed conditions. In terms of computational efficiency, the inference speed of both models reaches the millisecond level.